# PRISM: Proof-Carrying Artifact Generation through LLM $\times$ MDE Synergy and Stratified Constraints

### v1

@@ 3.3. Constraint Extraction and Integrated Constraint Model (ICM)
Finally, constraints admitted into the ICM must be turned into actual checkers used by CVG. Figure 5 illustrates PRISM’s stratified compilation pipeline.
Figure 5. Stratified compilation with LLM fallback and repository policy. Left: Channel 1 programmatic mapping compiles UMM/ICM constraints to L1 validators (JSON Schema, DFA, GBNF) and L2 validators (SHACL, SMT), preserving semantics. Right: when programmatic mapping is incomplete, an LLM-based pipeline (Ma et al., 2025b) synthesizes missing constraints with solver-backed validation; only verified outputs are persisted in the ICM storage.:
Following (Barrett et al., 2009; W3C RDF Data Shapes Working Group, 2017), we define compilation functions:
(14) $\kappa_{\mathrm{lex}}$ $:\mathcal{C}_{\mathrm{lex}}\rightarrow\mathrm{GBNF}$
(15) $\kappa_{\mathrm{struct}}$ $:\mathcal{C}_{\mathrm{struct}}\rightarrow\mathrm{DFA}_{\mathrm{prefix}}$
@@
Remark. In our implementation, context-free GBNF/CFG fragments are enforced at generation time using PDA/LR-style controllers. For bounded depth $d$, we apply finite unfolding to obtain an equivalent DFA for unified token masking and trace logging across structured decoding.
@@
Fix an unfolding bound $d$ for recursive structures. Let $\kappa_{\text{struct}}:\mathcal{C}_{\text{struct}}\!\to\!\text{DFA}$, $\kappa_{\text{sem}}:\mathcal{C}_{\text{sem}}\!\to\!\text{SHACL}$, $\kappa_{\text{logic}}:\mathcal{C}_{\text{logic}}\!\to\!\text{SMT}$, and write $\kappa$ for their union on $\mathcal{C}$ with cross-layer deduplication. Then:
@@ 3.4. Unified Automaton Execution with Theoretical Guarantees
Figure 6. Unified automaton execution for L1 enforcement during LLM decoding. Input constraints: L1 structural constraints ($\mathcal{C}_{\mathrm{struct}}$ from the ICM; Section 3.3) are compiled into executable automata. Finite-state fragments (JSON Schema / Regex / FSM) are determinized to prefix-closed DFAs. GBNF/CFG fragments are executed via PDA/LR; for unified masking and logging, bounded-depth unfolding (depth $d$) produces an equivalent DFA $\mathcal{A}_{d}$. Runtime decoding loop: At decoding step $t$, the executor computes $M_{t}=\{y\in\Sigma\mid\delta(s_{t},y)\neq\bot\}$, masks invalid tokens in the LLM logits, samples a valid token $y_{t}$, and transitions to $s_{t+1}=\delta(s_{t},y_{t})$. Audit trail: Each step records $\tau_{t}=\langle s_{t},|M_{t}|,y_{t},s_{t+1},\Delta t\rangle$ into the Evidence Registry. Post-generation, L2 validators (semantic checkers for graph constraints and logic checkers for temporal/numeric properties) certify the completed artifact, yielding composite evidence $\Pi=(\pi_{\text{struct}},\pi_{\text{sem}},\pi_{\text{logic}})\oplus\tau$ (Section 3.5).:
@@
Existing grammar-aligned decoding methods enforce a single global grammar or schema during LLM decoding and can align external grammars with subword vocabularies, using speculative decoding to keep runtime overhead negligible while preserving grammar validity (Scholak et al., 2021; Park et al., 2025a; Deljouyi, 2024; Beurer-Kellner et al., 2024; Park et al., 2025b; Lu et al., 2025). However, large-scale evaluations (e.g., SchemaBench / JSONSchemaBench) report two persistent failure modes: (i) over-constraining, where grammars block structurally unusual but semantically valid outputs, and (ii) under-constraining, where syntactically valid outputs still violate cross-reference integrity, safety rules, or timing constraints (Geng et al., 2025). The root cause is computability: not all constraints are prefix-decidable.
@@
This is not an engineering convenience but a theoretical split. L1 maximizes what can be proven safe online, and L2 certifies global semantics offline. The separation underpins PRISM’s claim of constraint-guided, auditable generation in safety-critical domains.
@@
Let $\mathcal{H}$ be a family of enforcement strategies, ordered by $\preceq$. For $h_{i},h_{j}\in\mathcal{H}$ we write $h_{i}\preceq h_{j}$ iff: (i) $\mathcal{L}(h_{i})\supseteq\mathcal{L}(h_{j})$, i.e., $h_{i}$ permits a superset of sequences accepted by $h_{j}$, and (ii) $\mathcal{P}(h_{i})\leq\mathcal{P}(h_{j})$, where $\mathcal{P}(\cdot)$ orders correctness guarantees. In PRISM:
@@
Enforcing DFA-based masking at L1 alters the model’s token distribution. This can bias the model toward minimally compliant completions, an effect observed in grammar-aligned decoding literature (Park et al., 2025a). We characterize this effect via parameterized grammars.
Definition 3.0 (Parameterized GBNF).
A parameterized GBNF is a tuple $\mathcal{G}_{\theta}=(N,\Sigma,R_{\theta},S)$ where $N$ is the set of non-terminals, $\Sigma$ is the token vocabulary, $S\in N$ is the start symbol, and $R_{\theta}$ is a set of production rules parameterized by $\theta\in\Theta$. Each production $A\to_{\theta}\alpha$ may include parameters $\theta_{k}$ that enable/disable alternatives or adjust rule preferences, where $\alpha\in(\Sigma\cup N\cup\{\theta_{k}\})^{*}$.
Execution semantics. GBNF/CFG fragments are enforced by a PDA/LR automaton maintaining a parse stack. When unified token masking and auditable traces are required, we apply bounded unfolding to depth $d$ to obtain an equivalent DFA $\mathcal{A}_{d}$ whose states encode stack configurations up to depth $d$. This preserves reachability guarantees (and thus prefix safety) within the bound $d$, while making the enforcement compatible with the DFA-style masking loop above.
Even with parameterized grammars, strict masking can distort generation quality. To mitigate this, PRISM employs three runtime strategies tailored to safety-critical model-driven engineering:
@@
    Minimum structural coverage. Before allowing EOS, the executor enforces configurable completeness thresholds: minimum token counts per structural component, presence of all semantically critical “optional” elements, and coverage of domain-specific fields. This prevents premature termination that would otherwise satisfy the DFA but yield under-specified artifacts.
@@
    Two-stage generation with targeted refinement. Phase 1 produces a structurally valid draft under tight DFA/PDA constraints (guaranteeing correctness). Phase 2 applies targeted refinement using AGR (Section 3.6), relaxing only semantic/logic aspects to enrich detail and style while preserving the structural invariants already certified by L1.
@@
    Calibrated unfolding depth. We tune the unfolding bound $d$ from Theorem 8 to preserve essential branching in $\mathcal{A}_{d}$. Shallower unfolding ($d<d_{\text{max}}$) reduces automaton size and runtime cost while still permitting diverse valid continuations, mitigating over-pruning and mode collapse.
In practice, these heuristics preserve usefulness and readability of generated artifacts while maintaining formal guarantees.
@@
Under the unified automaton execution framework, for any prefix $y_{1..t}$ produced by the verification-guided decoder over a prefix-closed DFA $\mathcal{A}$ (including bounded-unfolded $\mathcal{A}_{d}$), the current state $q_{t}$ satisfies: (i) $q_{t}\leadsto F$(there exists a continuation to acceptance); (ii) there exists a unique shortest completion $\mathcal{T}^{\star}(q_{t})$ of length $L(q_{t})=\min\{|w|:\delta^{*}(q_{t},w)\in F\}$; (iii) applying at most one structural edit based on $\mathcal{T}^{\star}(q_{t})$ yields an artifact that satisfies all L1 structural constraints.
@@
(i) and (ii) follow from the reachability-based masking policy and the breadth-first computation of $\mathcal{T}^{\star}(q_{t})$. For (iii), if the final state $q_{T}\notin F$ but $q_{T}\leadsto F$, appending $\mathcal{T}^{\star}(q_{T})$ performs a single closure step under the same structural parent in the parse tree. By construction this is the minimal edit that reaches $F$, bounding the structural repair parameter by $\lambda_{\text{struct}}\leq 1$. Empirically, we observe $\lambda_{\text{struct}}\approx 0.03$ (3% of cases require such repair) on AUTOSAR benchmarks (Section 4.3). ∎
Complexity Analysis.
Let $Q$ be the DFA state set after dead-state elimination, $k=|Q|$ the number of reachable states, and $|\Sigma|$ the token vocabulary size. Each decoding step examines at most $k\cdot|\Sigma|$ candidate transitions to derive $M_{t}$, yielding time complexity $\mathcal{O}(k\cdot|\Sigma|)$ per step and space complexity $\mathcal{O}(k)$ for state storage. The one-step structural closure (computing $\mathcal{T}^{\star}$ by breadth-first search) costs $\mathcal{O}(k)$. In practice, with $k\approx 10^{3}$ for AUTOSAR-scale schemas and $|\Sigma|\approx 10^{4}$ for modern LLM vocabularies, this overhead remains tractable (<1<1ms/step on commodity hardware).
@@ 3.5. Constraint-Guided Verifiable Generation
Role in PRISM.
Constraint-Guided Verifiable Generation (CVG) is the final stage of the PRISM pipeline. Given an artifact $a$ generated under Layer-1 (L1) structural guidance (Section 3.4) and validated under Layer-2 (L2) semantic/logic certification (Section 3.3), CVG assembles machine-checkable evidence that proves conformance to domain constraints and binds that evidence to the artifact. The result is a verifiable artifact that can be audited by an external authority without rerunning the (possibly proprietary) generation process. Unlike purely post-hoc verification approaches (Misu et al., 2024; ParthasarathyGaurav et al., 2024), PRISM records generation-time traces and post-generation validation results and fuses them into a single composite certificate.
@@
  • $a\in\mathcal{L}$ is the generated artifact (e.g., an AUTOSAR ARXML instance);
@@
Figure 7. PRISM’s validation and evidence composition. The generated artifact and the decode-time audit trail are checked by structural, semantic, and logic validators derived from the ICM. Each stage emits a machine-checkable trace. These traces are bundled into a single evidence package that is used both to accept or reject the artifact and to drive targeted repair when checks fail.:
@@
  • $s_{t}\in Q$ is the automaton configuration (DFA state or PDA stack) before emitting the token;
@@ 3.6. Audit-Guided Intelligent Repair
Role in PRISM.
By the time an artifact reaches Audit-Guided Repair (AGR), Layer-1 structural constraints have already been enforced. Prefix-closed DFA/PDA masking plus one-step completion (Theorem 12) guarantee that the draft artifact satisfies $\mathcal{C}_{\mathrm{struct}}$. Residual violations are therefore almost entirely Layer-2 failures: semantic graph obligations ($\mathcal{C}_{\mathrm{sem}}$, via SHACL) and logical/temporal/numeric obligations ($\mathcal{C}_{\mathrm{log}}$, via SMT). Traditional responses are either full regeneration (expensive, nondeterministic) or manual patching without guidance (error-prone). AGR replaces both with evidence-driven, localized repair.
@@
Both routes are driven by the same evidence $\Pi$, and both feed successful fixes back into PRISM as reusable constraints (“constraint promotion”), so future generations avoid repeating the violation. The two routes differ only in who performs the edit (LLM vs. human) and how the edit is applied (direct text patch vs. meta-model–aware configuration update). This “graduated automation” design is explicitly supported in the PRISM architecture: the Validation & Repair Layer and the Human Review Layer exchange artifacts, evidence, and compiled validators through the Evidence Registry and Configuration Editor.:contentReference[oaicite:1]index=1
@@
  • Semantic violations ($\pi_{\text{sem}}$). The SHACL validator returns, for each failing shape: (i) a violation path (an address into $a$ such as an AUTOSAR element or field), (ii) the expected condition (e.g., “must reference an existing Operation”), and (iii) a shape identifier. During ICM compilation (Section 3.3), each shape is annotated with ICM provenance: which ICM entry generated this rule, and which UMM entity it is anchored to. AGR uses this to (a) locate the exact broken reference in $a$, and (b) retrieve the authoritative rule and valid target types.
  • Logic violations ($\pi_{\text{logic}}$). The SMT validator produces an unsat-core $\mathcal{U}\subseteq\Phi$ that pinpoints the minimal inconsistent subset of numeric/temporal predicates. Each predicate in $\mathcal{U}$ is back-linked to concrete fields in $a$ and to its source obligation in the ICM (e.g., timing bounds, resource-usage limits). AGR can therefore highlight which values conflict and under which domain rule.
@@
In other words, AGR does not have to “search for the bug”: SHACL/SMT produce machine-readable failure tickets that name the failing node, the violated obligation, and its ICM/UMM provenance. This is crucial for both automated and human-assisted repair.
@@
Route A: Automated semantic/logic patching. AGR builds a repair prompt that includes: (i) the failing region of $a$, (ii) the violating constraint with natural-language explanation from ICM provenance, (iii) the expected target type(s) from the UMM, and (iv) any candidate values or references retrieved from the knowledge graph. The LLM (run at low temperature) proposes a minimal edit—for example, replacing an invalid cross-file reference with a valid one of the correct type, or reconciling two conflicting numeric bounds to satisfy the SMT model. The patched artifact $a^{\prime}$ is then locally re-validated on just that region to confirm the violation is cleared without introducing new ones.
Route B: Human-in-the-loop MDE editing. PRISM can also project the artifact $a$ into a configuration editor generated from Path S1 (Section 3.2). Path S1 converts domain XMI/XSD into a Unified Meta-Model (UMM) with typed entities, attributes, associations, and behavioral annotations (CRUD signatures, reference-lookup hooks). This UMM is compiled into strongly typed model classes plus a UI/editor that enforces field types, cardinalities, and reference pickers. In regulated domains (e.g., AUTOSAR), engineers already work with such meta-model-driven editors to author or adjust ARXML configurations.
AGR attaches validator diagnostics to that editor: for each SHACL/SMT violation, the Configuration Editor highlights the offending field(s), surfaces the failing rule (via ICM provenance), and offers valid candidates or safe ranges. A domain expert can then apply a compliant fix using familiar MDE-style interactions (selecting a valid referenced Operation from a dropdown, adjusting a timing bound within admissible limits), instead of free-text editing. After the human edit, the updated artifact is re-validated and re-materialized back into PRISM’s pipeline.:contentReference[oaicite:2]index=2
This route is not an afterthought: it is the fallback path in PRISM’s “graduated automation” model. Routine violations are auto-repaired, but high-stakes or ambiguous cases go through a configuration workflow that is already accepted by certification processes in safety-critical engineering: configure $\rightarrow$ validate $\rightarrow$ repair in an MDE tool.
@@
The repair convergence guarantee still holds under two observations: (i) Layer-1 structural conformance is already guaranteed and is not re-broken by AGR; and (ii) each remaining violation is either auto-fixable with probability >0.5>0.5 using ICM-grounded prompts, or is surfaced to a domain engineer through a configuration workflow that encodes the UMM and ICM constraints and is already standard practice in model-driven engineering toolchains (configure $\rightarrow$ validate $\rightarrow$ repair). This prevents blind retry loops and yields the empirical $\leq 1$-iteration repair convergence reported in Section 4.3.

### v2

@@ 3.3. Constraint Extraction and Integrated Constraint Model (ICM)
Finally, constraints admitted into the ICM must be turned into actual checkers used by CVG. Figure 5 illustrates ATLAS’s stratified compilation pipeline.
Figure 5. Stratified compilation with LLM fallback and repository policy. Left: Channel 1 programmatic mapping compiles UMM/ICM constraints to L1 validators (JSON Schema, DFA, GBNF) and L2 validators (SHACL, SMT), preserving semantics. Right: when programmatic mapping is incomplete, an LLM-based pipeline (maSpecGenAutomatedGeneration2025) synthesizes missing constraints with solver-backed validation; only verified outputs are persisted in the ICM storage.
Two-panel diagram of stratified compilation with LLM fallback and repository policy. Left panel: Channel 1 performs programmatic mapping from UMM/ICM constraints into L1, which enforces constraints at generation time using JSON Schema, FSM, GBNF, Regex, or DFA, and into L2, which enforces constraints post-generation via SHACL and SMT, while preserving intended semantics. Right panel: when the programmatic mapping is incomplete, a SpecGen-style pipeline invokes an LLM with solver-backed validation to synthesize missing SHACL/SMT constraints. Only constraints that pass validation are written into the Formal Constraints Database (ICM storage).
Following (barrett2009satisfiability; w3c2017shacl), we define compilation functions that map abstract constraints to executable enforcement mechanisms:
(12) $\kappa_{\mathrm{lex}}$ $:\mathcal{C}_{\mathrm{lex}}\rightarrow\mathcal{G}_{\mathrm{CFG}}$
(13) $\kappa_{\mathrm{struct}}$ $:\mathcal{C}_{\mathrm{struct}}\rightarrow\mathcal{A}_{\mathrm{prefix}}$
@@
MDE Interpretation: Compilation as Model Transformation.
Building upon the extraction methodology established in our S2D2 framework (DualStageFramework), which utilizes QVT-R to transform heterogeneous meta-models (XMI/XSD) into the UMM, ATLAS extends this chain by treating the compilation functions $\kappa$ as a subsequent set of vertical model transformations. Specifically, as visualized in Figure 6, $\kappa_{\mathrm{struct}}$ functions as a deterministic transformation rule that maps the static schema topology (from the ICM) into a dynamic state-transition system ($\mathcal{A}_{\mathrm{prefix}}$). Unlike the probabilistic nature of LLM token sampling, this transformation is mathematically rigorous and semantic-preserving (as supported by grammar-constrained decoding theory (scholakPICARDParsingIncrementally2021; poesiaSynchromeshReliableCode2021)). This distinction allows ATLAS to claim that Layer-1 enforcement is not merely a heuristic “guardrail,” but a correct-by-construction synthesis of the structural grammar, leaving only semantic ambiguities to the stochastic capabilities of the model.
Figure 6. Mapping the ATLAS constraint compilation pipeline to MDA concepts. The ICM functions as a Platform-Independent Model (PIM) of domain rules, which is deterministically transformed via $\kappa$ functions (Transformation Rules) into Platform-Specific Models (PSM) represented by L1 Prefix Automata and L2 Validators.
Remark. In our implementation, the context-free grammar $\mathcal{G}_{\mathrm{CFG}}$ is realized as GBNF/CFG fragments enforced via PDA/LR-style controllers. The prefix automaton $\mathcal{A}_{\mathrm{prefix}}$ is realized as a DFA (for regular constraints) or a bounded unfolding of the PDA (for recursive structures) to enable unified token masking.
@@
Fix an unfolding bound $d$ for recursive structures. Let $\kappa_{\text{struct}}$ map to prefix automata $\mathcal{A}_{\text{prefix}}$ (e.g., DFA), $\kappa_{\text{sem}}$ map to semantic validators $\mathcal{V}_{\text{sem}}$ (e.g., SHACL), and $\kappa_{\text{logic}}$ map to logic validators $\mathcal{V}_{\text{log}}$ (e.g., SMT). Write $\kappa$ for their union on $\mathcal{C}$ with cross-layer deduplication. Then:
@@ 3.4. Unified Automaton Execution with Theoretical Guarantees
Figure 7. Unified automaton execution for L1 enforcement during LLM decoding. Input constraints: L1 structural constraints ($\mathcal{C}_{\mathrm{struct}}$ from the ICM; Section 3.3) are compiled into executable automata. Finite-state fragments (JSON Schema / Regex / FSM) are determinized to prefix-closed DFAs. GBNF/CFG fragments are executed via PDA/LR; for unified masking and logging, bounded-depth unfolding (depth $d$) produces an equivalent DFA $\mathcal{A}_{d}$. Runtime decoding loop: At decoding step $t$, the executor computes $M_{t}=\{y\in\Sigma\mid\delta(s_{t},y)\neq\bot\}$, masks invalid tokens in the LLM logits, samples a valid token $y_{t}$, and transitions to $s_{t+1}=\delta(s_{t},y_{t})$. Audit trail: Each step records $\tau_{t}=\langle s_{t},|M_{t}|,y_{t},s_{t+1},\Delta t\rangle$ into the Evidence Registry. Post-generation, L2 validators (semantic checkers for graph constraints and logic checkers for temporal/numeric properties) certify the completed artifact, yielding composite evidence $\Pi=(\pi_{\text{struct}},\pi_{\text{sem}},\pi_{\text{logic}})\oplus\tau$ (Section 3.5).
Diagram of unified automaton execution for Layer 1 (L1) enforcement during large language model decoding. First, structural constraints from the Integrated Constraint Model (ICM) are compiled into executable automata: JSON Schema, Regex, and FSM fragments become deterministic finite automata, while GBNF and CFG fragments are executed via PDA/LR and optionally unfolded to bounded-depth DFAs for unified masking and logging. During decoding at each step t, the executor computes the set of valid next tokens, masks invalid tokens in the model logits, samples a valid token, and advances the automaton state. Each decoding step logs an audit tuple including states, mask size, chosen token, and timing. After generation finishes, Layer 2 (L2) validators for semantic constraints (e.g., SHACL) and logical constraints (e.g., SMT) and SMT (logical and temporal constraints) certify the final artifact and combine the audit trail into composite evidence.
@@
Existing grammar-aligned decoding methods enforce a single global grammar or schema during LLM decoding and can align external grammars with subword vocabularies, using speculative decoding to keep runtime overhead negligible while preserving grammar validity (scholakPICARDParsingIncrementally2021; parkGrammaralignedDecoding2025; deljouyiUnderstandableTestGeneration2024; zotero-item-1597; Park2025; Lu2025b). However, large-scale evaluations (e.g., SchemaBench / JSONSchemaBench) report two persistent failure modes: (i) over-constraining, where grammars block structurally unusual but semantically valid outputs, and (ii) under-constraining, where syntactically valid outputs still violate cross-reference integrity, safety rules, or timing constraints (zotero-item-1624). The root cause is computability: not all constraints are prefix-decidable.
@@
This is not an engineering convenience but a theoretical split. L1 maximizes what can be proven safe online, and L2 certifies global semantics offline. The separation underpins ATLAS’s claim of constraint-guided, auditable generation in safety-critical domains.
@@
Let $\mathcal{H}$ be a family of enforcement strategies, ordered by $\preceq$. For $h_{i},h_{j}\in\mathcal{H}$ we write $h_{i}\preceq h_{j}$ iff: (i) $\mathcal{L}(h_{i})\supseteq\mathcal{L}(h_{j})$, i.e., $h_{i}$ permits a superset of sequences accepted by $h_{j}$, and (ii) $\mathcal{P}(h_{i})\leq\mathcal{P}(h_{j})$, where $\mathcal{P}(\cdot)$ orders correctness guarantees. In ATLAS:
@@
Enforcing DFA-based masking at L1 alters the model’s token distribution, effectively pruning the probability mass of structurally invalid tokens to zero. While ensuring correctness, this strict truncation can exacerbate distribution distortion: the model may be biased toward the shortest valid path (e.g., skipping optional fields or closing lists early) to maximize per-step probability, resulting in syntactically correct but semantically impoverished artifacts (parkGrammaralignedDecoding2025).
To mitigate this, ATLAS employs three runtime strategies tailored to safety-critical model-driven engineering, applicable across both DFA and PDA execution modes:
@@
    Minimum structural coverage. Instead of purely relying on the automaton’s accepting states ($F$) to allow termination, the executor enforces configurable completeness thresholds. For example, specific transitions leading to the End-of-Sequence (EOS) token are masked until minimum element counts are met or mandatory ”optional” fields (e.g., unique IDs) are generated. This forces the model to explore deeper structural paths.
@@
    Two-stage generation with targeted refinement. Phase 1 produces a structurally valid draft under tight constraints. Phase 2 applies targeted refinement using AGR (Section 3.6), relaxing only semantic/logic aspects to enrich detail while preserving the structural invariants already certified by L1.
@@
    Calibrated unfolding depth. We tune the unfolding bound $d$ from Theorem 7. Shallower unfolding reduces automaton size while still permitting diverse valid continuations, preventing the ”over-pruning” that often leads to mode collapse in deeply nested recursive grammars.
In practice, these heuristics balance the rigorous requirements of formal verification with the linguistic diversity of LLMs.
@@
Under the unified automaton execution framework, for any prefix $y_{1..t}$ produced by the verification-guided decoder over a prefix-closed DFA $\mathcal{A}$ (including bounded-unfolded $\mathcal{A}_{d}$), the current state $q_{t}$ satisfies: (i) $q_{t}\leadsto F$(reachability to acceptance is maintained); (ii) there exists a unique shortest completion suffix $\mathcal{T}^{\star}(q_{t})$; (iii) applying a single closure operation (appending $\mathcal{T}^{\star}(q_{t})$) yields an artifact that satisfies all L1 structural constraints.
@@
(i) and (ii) follow from the reachability-based masking policy. For (iii), if generation halts at state $q_{t}\notin F$, the masking guarantee $q_{t}\leadsto F$ ensures a path exists. $\mathcal{T}^{\star}(q_{t})$ represents the minimal sequence of tokens required to traverse from $q_{t}$ to the nearest state in $F$. Appending this suffix constitutes a single deterministic repair action ($\lambda_{\text{struct}}\leq 1$), ensuring the final artifact is structurally valid by construction. ∎
Runtime Overhead.
The unified automaton execution imposes negligible latency. Since the allowed token set $M_{t}$ depends only on the current state’s outgoing transitions, the per-step masking overhead is minimal (<1<1ms on commodity hardware), scaling efficiently with the vocabulary size $|\Sigma|$ and the complexity of the constraint schema.
@@ 3.5. Constraint-Guided Verifiable Generation
Role in ATLAS.
Constraint-Guided Verifiable Generation (CVG) is the final stage of the ATLAS pipeline. Given an artifact $a$ generated under Layer-1 (L1) structural guidance (Section 3.4) and validated under Layer-2 (L2) semantic/logic certification (Section 3.3), CVG assembles machine-checkable evidence that proves conformance to domain constraints and binds that evidence to the artifact. The result is a verifiable artifact that can be audited by an external authority without rerunning the (possibly proprietary) generation process. Unlike purely post-hoc verification approaches (misuAIAssistedSynthesisVerified2024; parthasarathygauravTrustworthyAutomatedProgram2024), ATLAS records generation-time traces and post-generation validation results and fuses them into a single composite certificate.
@@
  • $a\in\mathcal{L}$ is the generated artifact, formally a structured instance model conforming to the M2-level UMM (realized as ARXML, JSON, or source code AST depending on the domain);
@@
Figure 8. ATLAS’s validation and evidence composition. The generated artifact and the decode-time audit trail are checked by structural, semantic, and logic validators derived from the ICM. Each stage emits a machine-checkable trace. These traces are bundled into a single evidence package that is used both to accept or reject the artifact and to drive targeted repair when checks fail.
Workflow for validation and evidence composition in ATLAS. Inputs include the generated artifact, the runtime audit trail from Layer 1 guided decoding, and validators compiled from the Integrated Constraint Model (ICM): Layer 1 structural validators such as DFA, JSON Schema, and GBNF, and Layer 2 semantic and logical validators such as SHACL and SMT, including constraints produced with LLM assistance. Validation executes in structural, then semantic, then logic order. Each stage produces a machine-checkable trace that is appended to an Evidence Registry. These traces are combined into a composite evidence package Pi together with the audit trail. A Trace Verifier evaluates this package; if it passes, the artifact plus evidence is sealed and exposed as a verifiable artifact, and if it fails, an Audit-Guided Repair process uses the evidence to generate targeted fixes.
@@
  • $s_{t}$ is the automaton configuration (state $q\in Q$ for DFA or stack configuration for PDA) before emitting the token, aligning with the execution semantics in Section 3.4;
@@ 3.6. Audit-Guided Intelligent Repair
Role in ATLAS.
By the time an artifact reaches Audit-Guided Repair (AGR), Layer-1 structural constraints have already been enforced. Prefix-closed DFA/PDA masking plus one-step completion (Theorem 10) guarantee that the draft artifact satisfies $\mathcal{C}_{\mathrm{struct}}$. Residual violations are therefore almost entirely Layer-2 failures: semantic graph obligations ($\mathcal{C}_{\mathrm{sem}}$) and logical/temporal/numeric obligations ($\mathcal{C}_{\mathrm{log}}$). While recent iterative refinement strategies leverage feedback loops (e.g., compiler logs or test outputs) (shinn2023reflexion; chen2024selfdebug), they often lack the formal provenance required to link failures back to normative standards. AGR advances this by employing an evidence-driven, localized repair strategy backed by formal or verifiable programmatic validators.
@@
Both routes are driven by the same evidence $\Pi$, and both feed successful fixes back into ATLAS as reusable constraints (“constraint promotion”), so future generations avoid repeating the violation. The two routes differ only in who performs the edit (LLM vs. human) and how the edit is applied (direct text patch vs. meta-model–aware configuration update). This “graduated automation” design is explicitly supported in the ATLAS architecture: the Validation & Repair Layer and the Human Review Layer exchange artifacts, evidence, and compiled validators through the Evidence Registry and Configuration Editor.:contentReference[oaicite:1]index=1
@@
  • Semantic violations ($\pi_{\text{sem}}$). The semantic validator (instantiated formally or programmatically) returns, for each violation: (i) a violation path (an address into $a$ such as an entity ID or XPath), (ii) the expected condition (e.g., “must reference an existing Operation”), and (iii) a constraint identifier. During ICM compilation (Section 3.3), each constraint is annotated with ICM provenance: which ICM entry generated this rule, and which UMM entity it is anchored to. AGR uses this to (a) locate the exact locus of failure in $a$, and (b) retrieve the authoritative rule and valid target types from the UMM.
  • Logic violations ($\pi_{\text{logic}}$). The logic validator identifies a minimal conflict set (e.g., an SMT unsat-core or a failed assertion trace) $\mathcal{U}\subseteq\Phi$ that pinpoints the inconsistent subset of numeric/temporal predicates. Each predicate in $\mathcal{U}$ is back-linked to concrete fields in $a$ and to its source obligation in the ICM (e.g., timing bounds, resource-usage limits). AGR can therefore highlight which values conflict and under which domain rule.
@@
In other words, AGR does not have to “search for the bug”: the L2 validators produce machine-readable failure tickets that name the failing node, the violated obligation, and its ICM/UMM provenance. This is crucial for both automated and human-assisted repair.
@@
Route A: Automated validation-guided regeneration. Unlike blind retry strategies, AGR utilizes the violation reports (e.g., SHACL violation paths or SMT unsat cores) to construct a constraint-aware repair prompt. This prompt explicitly injects the violated rules and the relevant UMM schema context, guiding the LLM to regenerate the specific artifact file (or affected block) with corrected logic. This represents a diagnostic-driven regeneration at the file granularity, which effectively narrows the search space without requiring token-level surgical editing.
Route B: Human-in-the-loop MDE editing. ATLAS acknowledges that for complex, global semantic violations (the “last 20%”), fully automated repair may not converge efficiently. In these cases, ATLAS projects the artifact into a context-aware configuration editor generated via Path S1. Crucially, AGR attaches the machine-readable validator diagnostics directly to the editor’s UI. This transforms the human role from manual authoring to evidence-based decision making, resolving ambiguities that are computationally undecidable for the LLM while leveraging the UMM structure to ensure valid edits.
This route is not an afterthought: it is the fallback path in ATLAS’s “graduated automation” model. Routine violations are auto-repaired, but high-stakes or ambiguous cases go through a configuration workflow that is already accepted by certification processes in safety-critical engineering: configure $\rightarrow$ validate $\rightarrow$ repair in an MDE tool.
@@
The repair convergence guarantee still holds under two observations: (i) Layer-1 structural conformance is already guaranteed and is not re-broken by AGR; and (ii) each remaining violation is either auto-fixable with probability >0.5>0.5 using ICM-grounded prompts, or is surfaced to a domain engineer through a configuration workflow that encodes the UMM and ICM constraints and is already standard practice in model-driven engineering toolchains (configure $\rightarrow$ validate $\rightarrow$ repair).